Circular No. 6178

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

MARKARIAN 501 The Whipple Observatory Gamma Ray Collaboration (J. Quinn, J.

Buckley, and T. C. Weekes, Smithsonian Astrophysical Observatory; M. Catanese, D. A. Carter-Lewis, F. Krennrich, R. C. Lamb, and J. Zweerink, Iowa State University; C. W. Akerlof, D. Meyer, and M. Schubnell, University of Michigan; J. Gaidos, G. Sembroski, and C. Wilson, Purdue University; D. J. Fegan, R. Lessard, and J. McEnery, University College, Dublin; M. F. Cawley, St. Patrick’s College, Maynooth; and S. Biller, A. M. Hillas, A. Rodgers, and J. Rose, University of Leeds) communicates: “TeV γ-rays have been detected from the BL Lac-object Markarian 501 with the highresolution imaging camera on the Whipple 10-m optical reflector between Mar. 28 and June 5. A flux was detected in 57 hr of observation at the 10σ level. The flux above 0.3 TeV was ∼ 8 × 10^{−12} photon cm^{−1} s^{−1} (or 0.07 that of the Crab Nebula). The TeV-γ-ray source’s location is within 10′ of the known position of Mrk 501 (α = 16^{h}52^{m}.2, δ = +39^{o}50′, equinox 1950.0) and consistent with this identification. Mrk 501, at a redshift of 0.034, is similar in some respects to the previously detected TeV-BL Lac source Mrk 421 (Kerrick et al. 1995, Ap.J. 438, L59, and references therein); however, it does not yet show the same evidence for TeV variability as Mrk 421, nor has it yet been detected at MeV to GeV γ-ray energies.”

BL LACERTAE R. D. Jeffries and M. R. Thurston, University of Birmingham; and

D. J. Axon, Space Telescope Science Institute, communicate: “We have observed BL Lac in photometric conditions with the 1-m Jacobus Kapetyn telescope (+ CCD camera) at La Palma. A comparison with a photometric sequence in the CCD field yields: June 4.21 UT, V = 15.76 ± 0.02, B − V = +1.03 ± 0.02; 6.18, 15.61 ± 0.02, 0.97 ± 0.03; 7.20, 15.51 ± 0.02, 0.97 ± 0.02, U − B= −0.16 ± 0.03. This confirms the suspicion of Vermeuelen et al. (IAUC 6176) that BL Lac is currently in a very low state at around its historical minimum.”

COMET 71P/CLARK Visual magnitude estimates by A. Hale, Cloudcroft, NM (0.41-m re-

flector): May 28.44 UT, 11.3; June 3.40, 11.2.

1995 June 16 Daniel W. E. Green